OpenStudy (anonymous):

How do you find the polynomial function that has the given zeroes? 0, -2, -3

OpenStudy (anonymous):

x*(x+2)*(x+3)

jimthompson5910 (jim_thompson5910):

x = 0, x = -2, or x = -3 x = 0, x + 2 = 0 or x+3 = 0 x(x+2)(x+3)=0 x(x^2+5x+6)=0 x^3+5x^2+6x=0 Polynomial function is f(x)=x^3+5x^2+6x

OpenStudy (anonymous):

X^3+5x^2+6x=0